Please consider the following values for the variables X and Y. Treat each row as a pair of scores for the variables X and Y (with the first row providing the labels "X" and "Y").

X Y
2 4
4 3
6 5
7 7
11 6

Please calculate Pearson's correlation coefficient (r) for these data and report your answer below. When reporting your answer, please provide three decimal places (if relevant).
Answer #1

Solution:

x y x2 y2 x⋅y
2 4 4 16 8
4 3 16 9 12
6 5 36 25 30
7 7 49 49 49
11 6 121 36 66
x=30 y=25 x2=226 y2=135 xy=165

r = n ( xy )- (x ) (y) / n [ ( x2 ) - ( x )2 ] [ n (y2 ) - ( y )2

r = 5 (165) - (30 * 25 )/ [ 5 * 226 - ( 30 ) 2   [ 5 * 135 - ( 25 ) 2 ]

r = 825 - 750 / 1130 - 900  675 -625

r = 75 /  23050

r = 75 / 15.1658 * 7.0711

r = 75 / 107.2381

r = 0.699